# Semantic Transmission Framework in Direct Satellite Communications

### v1

@@ I Introduction
With the rapid development of sixth-generation (6G), satellite communications have become an important avenue for future wireless networks due to their capability to provide seamless connectivity to global ground users [1]. Compared with traditional terrestrial communication systems, satellite communications can effectively overcome geographical limitations and guarantee service availability without ground infrastructure. Thus, it is particularly suitable for emergency communication scenarios. In recent years, the rapid deployment of satellite constellations such as Starlink and OneWeb, have significantly accelerated the construction of large-scale satellite communication networks, which opened new frontiers for applications in cognitive communications [2], cell-free networks [3], and edge computing [4].
However, due to limited transmit power, constrained bandwidth, and extremely long transmission distances, direct satellite communications face stringent link budget limitations [5]. Long distance transmission between satellites and ground communication devices suffers from severe signal attenuation and propagation delay, spectrum efficiency becomes increasingly important for future satellite networks. To address this challenge, semantic communications, which aim to enable meaning-oriented information transmission, have emerged for future wireless communications in recent years [6]. By understanding the semantics of the transmitted data, semantic communications extract semantic features and transmit them to the receiver, then the receiver can reconstruct the data based on these features. As a result, semantic communications can significantly reduce the required radio resources to enhance the spectral efficiency, rendering it particularly suitable for direct satellite communications under low signal-to-noise ratio (SNR) conditions [7].
@@ III Decision-Assisted DRL-Based Optimization
To be specific, we apply action masking on the policy logits: for any $a\notin\mathcal{A}_{\rm feas}(s_{t})$, its logit is set to -$\infty$ before the softmax so that $\pi_{\theta}(a|s_{t})=0$, and the remaining probabilities are renormalized. This method narrows the range of the original large action space into a smaller feasible one, therefore, this method can significantly reduce the variance in policy-gradient updates.

### v2

# arXiv:2601.00381v2[cs.IT] 27 Feb 2026

@@ I Introduction
With the rapid development of sixth-generation (6G), satellite communications have become an important avenue for future wireless networks due to their capability to provide seamless connectivity to global ground users [1]. Compared with terrestrial communication systems, satellite communications can effectively overcome geographical limitations and guarantee service availability without ground infrastructure. Thus, it is particularly suitable for emergency communication scenarios. In recent years, the rapid deployment of satellite constellations such as Starlink and OneWeb, have significantly accelerated the construction of large-scale satellite communication networks and bring new frontiers for applications in cognitive communications [2], rate splitting multiple access (RSMA) [3], and edge computing [4].
However, due to limited transmit power, constrained bandwidth, and extremely long transmission distances, direct satellite communications face stringent link budget limitations [5]. Long distance transmission between satellites and ground communication devices suffers from severe signal attenuation and propagation delay, spectrum efficiency becomes increasingly important for future satellite networks. To address this challenge, semantic communications, which aim to enable meaning-oriented information transmission, have emerged for future wireless communications in recent years [6]. By understanding the semantics of the data, semantic communications extract semantic features and transmit them to the receiver, then the receiver can reconstruct the data based on these features. As a result, semantic communications can significantly reduce the required radio resources to enhance the spectral efficiency, rendering it particularly suitable for satellite communications under limited spectrum resource conditions [7].
@@ III Decision-Assisted DRL-Based Optimization
To be specific, we apply action masking on the policy logits: for any $a\notin\mathcal{A}_{\rm feas}(s_{t})$, its logit is set to -$\infty$ before the softmax so that $\pi_{\theta}(a|s_{t})=0$, and the remaining probabilities are renormalized. This method narrows the range of the original large action space into a smaller feasible one, therefore, this method can significantly reduce the variance in policy-gradient updates. Notice that due to the space limitation, other technical details can refer to the original algorithm paper [17].